中国分布式计算论坛

 找回密码
 新注册用户
搜索
楼主: vmzy

[BOINC] [天文类] SETI@home

[复制链接]
发表于 2005-12-24 12:37:20 | 显示全部楼层

December 23, 2005

Yesterday we stopped accepting results for SETI@home Classic. The Classic project stats are now frozen.  

昨天SETI@home Classic已经停止接收计算结果了。Classic的项目统计也已冻结。

评分

参与人数 1基本分 +10 维基拼图 +3 收起 理由
BiscuiT + 10 + 3

查看全部评分

发表于 2005-12-25 11:01:08 | 显示全部楼层
引用 Youth 在 2005-12-24 12:37 时的帖子:
Yesterday we stopped accepting results for SETI@home Classic. The Classic project stats are now frozen.  

昨天SETI@home Classic已经停止接收计算结果了。Classic的项目统计也已冻结。


结束了~叹息~各位算seti c 的前辈何去何从啊?
发表于 2005-12-25 11:06:04 | 显示全部楼层
估计有些“收工”了,有些转去别的项目了吧~

[ Last edited by Youth on 2006-2-14 at 16:11 ]
发表于 2005-12-28 09:03:04 | 显示全部楼层

December 27, 2005

We have temporarily turned off the counting of workunits and results in various states (in progress, waiting for validation, etc) in order to give the database cleanup process more resources. Until we turn the counts back on, several of the numbers on the server status page will not be up to date.

我们临时关闭了任务包及计算结果的状态计数(比如计算中,等待验证等等),以便给数据库清理进程更多的处理能力。服务器状态页面上的数据将不再更新,直到我们重新启动计数。

评分

参与人数 1基本分 +10 维基拼图 +5 收起 理由
BiscuiT + 10 + 5

查看全部评分

 楼主| 发表于 2006-1-6 14:30:21 | 显示全部楼层
January 5, 2006 - 23:30 UTC
We're back on line after a day-and-a-half long outage during which we were able to clear out enough of our BOINC database to get us back to the same levels as in early December. Of course, there's going to be a painful period as all the BOINC clients clog our servers with demands for work. This will push through eventually as it always has. Meanwhile the master database merge is going along swimmingly - the result and signals tables are copying over much faster than expected.

2006年1月6日 - 7:30 北京时
在一天半的停机后,我们恢复正常了,在停机期间里,我们为BOINC数据库清理出了和12月上旬的水平相当的足够的空间。当然,那是一段痛苦的时间,因为所有BOINC客户端都在通过要求新工作,来防碍我们的服务器。不过它像往常一样最终艰难的完工了。同时主数据库合并得以顺利进行 - 结果和信号表的复制速度比期望的要快。

评分

参与人数 1基本分 +20 维基拼图 +8 收起 理由
BiscuiT + 20 + 8

查看全部评分

 楼主| 发表于 2006-1-8 00:22:42 | 显示全部楼层
January 5, 2006
The final SETI@home Classic work totals have been copied to SETI@home/BOINC, and are shown on User Pages.

2006年1月5日
SETI@home Classic的任务信息终于全部拷到SETI@home/BOINC里了,显示在User页面里!

评分

参与人数 1基本分 +10 维基拼图 +5 收起 理由
BiscuiT + 10 + 5

查看全部评分

发表于 2006-1-8 11:51:53 | 显示全部楼层

January 7, 2006

There is a lab-wide problem (not just SETI@home) that is causing our connection to the world to briefly disappear every few minutes. Otherwise everything is working just fine.

实验室对外的网络出了问题,会时不时地断开。其它一切正常。

评分

参与人数 1基本分 +10 维基拼图 +3 收起 理由
BiscuiT + 10 + 3

查看全部评分

 楼主| 发表于 2006-1-10 22:47:04 | 显示全部楼层
January 9, 2006
Today (Monday) at 18:00 UTC we will have a short outage to compress and back up our database. The data servers and most web pages will be offline during this time. The outage will last 3-4 hours, and then there is a long period of "catch up" afterwards as the servers are swamped with backlogged requests.

2006年1月9日
今天(星期一)在18:00 UTC,我们将有短暂的停机,以压缩和备份我们的数据库。此间数据服务器和多数网页将离线。停机将持续3-4个小时,并且然后会有很长一段时间服务器才能"真正恢复",因为会陷入大量的请求响应之中。

评分

参与人数 1基本分 +10 维基拼图 +6 收起 理由
BiscuiT + 10 + 6

查看全部评分

发表于 2006-1-15 11:21:40 | 显示全部楼层

January 13, 2006

There have been numerous but very brief network drop outs over the past few days affecting all of the Space Sciences Lab. This turned out to be a problem with the SSL firewall and is hopefully fixed at this point.

过去几天网络有问题,影响了整个实验室。最终发现是SSL防火墙的问题,很快将进行修复。

评分

参与人数 1基本分 +10 维基拼图 +3 收起 理由
BiscuiT + 10 + 3

查看全部评分

发表于 2006-1-15 11:24:06 | 显示全部楼层

January 14, 2006

SETI@home Classic users: if you've been unable to link your BOINC and Classic accounts (with the error message 'No SETI@home/BOINC account with Classic ID xxxx was found') please try again; we fixed a bug in the script.

致Classic的用户:如果你们不能将SETI/BOINC帐号和Classic帐号关联(错误信息是“SETI@home/BOINC account with Classic ID xxxx was found'”),请重新试一次,我们刚修复了脚本中的一个错误。

评分

参与人数 1基本分 +10 维基拼图 +5 收起 理由
BiscuiT + 10 + 5

查看全部评分

跨越地平线 该用户已被删除
发表于 2006-1-15 11:33:27 | 显示全部楼层
提示: 作者被禁止或删除 内容自动屏蔽
发表于 2006-1-17 09:33:42 | 显示全部楼层

January 16, 2006

It looks like we may be experiencing local SSL network problems again. If so, this will likely be fixed tomorrow. Today is a university holiday. In the meantime there will be intermittent problems contacting the scheduling server.

SSL网络的问题似乎又出现了,明天应该能修复,今天是学校的假日。期间大家连接调度服务器时可能会有问题。

评分

参与人数 1基本分 +10 维基拼图 +4 收起 理由
BiscuiT + 10 + 4

查看全部评分

发表于 2006-1-20 10:52:32 | 显示全部楼层

估计是在2006年1月16日~17日发出的

No Work Notice
We are almost done with the science database merge. During this operation we are unable to create new work, so we eventually ran out of data to send to our participants. We hope finish up and start creating new work later this evening or by tomorrow morning.
!断粮通告!
数据库合并基本完成。在此期间我们不能制造新的wu,所以现有的wu几乎已经派发完了。我们希望尽量能在今天晚上或明天早上给出新的wu。

[ Last edited by Youth on 2006-2-14 at 16:11 ]

评分

参与人数 1基本分 +10 维基拼图 +5 收起 理由
BiscuiT + 10 + 5

查看全部评分

发表于 2006-1-20 11:04:51 | 显示全部楼层
断粮通告:今天晚上或明天早上给出新的WU!!
我就说,今,算完了1个,只上传了结果,没能下到新任务呀……
2006-1-20 10:43:21|SETI@home|Scheduler request to http://setiboinc.ssl.berkeley.edu/sah_cgi/cgi failed
2006-1-20 10:43:21|SETI@home|No schedulers responded
2006-1-20 10:43:22|SETI@home|Deferring communication with project for 58 seconds
还是楼上的消息灵通呀!

[ Last edited by Youth on 2006-2-14 at 16:11 ]
发表于 2006-1-20 12:05:18 | 显示全部楼层

17日的技术新闻

January 17, 2006 - 23:30 UTC

The final big stage of the database merge started this morning, which is why the assimilators and splitters are off. Basically, we are now merging the final bits of data that were collected in the last month (since we started the merge). This should take a couple of days - hopefully we have enough work to send to last that long.
Over the long weekend we had a couple of connectivity issues. Both the upload and download servers have some kind of automounter problem where, under heavy load, they suddenly lose a random subset of their mounts. When this happens they still function but in a somewhat degraded mode, and eventually need to be rebooted. We also still have that feeder bug which clogs its shared memory segment so it is unable to cache enough work to send out. Both of these problems are being diagnosed. There are also continuing issues with the lab firewall/routers which is being handled by the lab administrators. Things are much better now in general, but there may be a day or two of the usual "catch up".

On the plus side, since we turned off the classic data server, we freed up one of our precious few Cogent IP addresses. So we are working toward moving the scheduler onto our private Cogent link, which will keep the bandwidth off the lab routers and therefore protected from those problems.

评分

参与人数 1基本分 +5 收起 理由
BiscuiT + 5

查看全部评分

您需要登录后才可以回帖 登录 | 新注册用户

本版积分规则

论坛官方淘宝店开业啦~

小黑屋|手机版|Archiver|中国分布式计算总站 ( 沪ICP备05042587号 )

GMT+8, 2022-8-19 00:25

Powered by Discuz! X3.4

© 2001-2017 Comsenz Inc.

快速回复 返回顶部 返回列表